OpenStudy (anonymous):

medal for answer Thank you The binomial expansion of (2x-y)^3 is

OpenStudy (manish00333):

try finding it on your own! it's just: \[\left( 2x-y\right)\left( 2x-y\right)\left( 2x-y\right)\]

OpenStudy (anonymous):

so is it 8x^3-12x^2y+6xy^2-y^3?

OpenStudy (anonymous):

Thats what i think it is but I'm not too sure

OpenStudy (manish00333):

yep, that's correct :)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

thanks for helping out!

OpenStudy (manish00333):

np :P